# Personalize to generalize: Towards a universal medical multi-modality generalization through personalization

### v1

@@ Abstract
Personalized medicine is a groundbreaking healthcare framework for the $21^{st}$ century, tailoring medical treatments to individuals based on unique clinical characteristics, including diverse medical imaging modalities. Given the significant differences among these modalities due to distinct underlying imaging principles, generalization in multi-modal medical image tasks becomes substantially challenging. Previous methods addressing multi-modal generalization rarely consider personalization, primarily focusing on common anatomical information. This paper aims to bridge multi-modal generalization with the concept of personalized medicine. Specifically, we propose a novel approach to derive a tractable form of the underlying personalized invariant representation $\mathbb{X}_{h}$ by leveraging individual-level constraints and a learnable biological prior. We demonstrate the feasibility and benefits of learning a personalized $\mathbb{X}_{h}$, showing that this representation is highly generalizable and transferable across various multi-modal medical tasks. Our method is rigorously validated on medical imaging modalities emphasizing both physical structure and functional information, encompassing a range of tasks that require generalization. Extensive experimental results consistently show that our approach significantly improves performance across diverse scenarios, confirming its effectiveness.
@@ 1 Introduction
Figure 1: A diagram of medical modalities and individual differences. Individual variations may be significant and warrant further research attention from the medical intelligence community.
Personalized medicine represents a transformative framework for $21^{st}$ century healthcare, tailoring medical treatments to each patient’s unique characteristics [45, 24, 4]. This approach necessitates diverse information, including clinical data such as radiological images. Three-dimensional medical images, generated through specialized techniques and radiopharmaceuticals, excel at highlighting specific anatomical features. Collectively, different medical image modalities provide a comprehensive view of a patient’s structural and functional characteristics. However, this distinction between medical modalities creates significant generalization challenges in medical image analysis, especially when certain modalities may be inaccessible due to an individual’s financial constraints or physical limitations, thereby complicating the effectiveness of personalized medicine.
As illustrated in Fig. 1, contemporary research in medical intelligence is mainly concentrated on structural modalities that depict physical anatomy. This includes Magnetic Resonance Imaging (MRI) scans [53], which use strong magnetic fields and radiofrequency currents yielding distinct sequences, and Computed Tomography (CT) scans [33, 51], which employ X-rays to measure its attenuation. Other studies [49] focus on the functional modalities associated with biochemistry, such as Positron Emission Tomography (PET) scans. PET scans are expensive functional imaging scans that employ radiotracers emitting gamma rays to visualize and measure metabolic processes. The differing imaging principles result in substantial modality gaps, presenting a critical challenge for model generalization. For clarity, we categorize these modalities for generalization tasks into two types: homogeneous generalization, which pertains to generalizing within structural or functional modalities (e.g., T1, T2, T1ce, and Flair in MRI, as shown in Fig. 1); and heterogeneous generalization, which involves generalizing across both structural and functional modalities, such as CT and PET.
As each person is fundamentally different from the average of the population [45], the concept of multi-modality generalization needs comprehensive discussion within the scope of personalization, an area scarcely addressed in previous research. An ideally well-generalized personalized medical model across modalities should (1) provide additional insights derived from all modalities when only a subset is accessible and (2) seamlessly transfer across domains while maintaining the capacity to achieve (1). While pre-training can significantly enhance downstream generalization, some recent approaches concentrate on learning common physical anatomy invariance at the class level [23], which may overlook the individual variations. Another line of research [40, 46, 23] primarily addresses the transferability of single-modal tasks and may not be suitable for multi-modal scenarios. In addition, most research on homogeneous generalization for medical tasks focuses on structural sequences of MRI or CT, employing strategies such as cross-modality transfer [30, 25, 51] or targeting challenges like missing modality segmentation [31, 6, 36, 37, 51]. Heterogeneous generalization presents a greater challenge due to the disparities between structural and functional information. Despite its significance, very few efforts [34] have addressed heterogeneous generalization, mainly focusing on one-directional modality transfer (e.g., PET to CT or MRI), and rarely exploring the model’s generalizability and transferability for other tasks in this context.
To address multi-modality generalization issue for personalization under both homogeneous and heterogeneous settings, we formally introduce the concept of the personalized invariant representation for multi-modal generalization, denoted as $\mathbb{X}_{h}$, and its constraints as outlined in Hypothesis 3.1. Furthermore, personalized invariant $\mathbb{X}_{h}$, which learns aggregated biological information from all possible modalities specific to the individual, is likely to enhance performance across various medical tasks for that person. Building on this hypothesis, this paper proposes a general approach aimed at enhancing the generalization of various medical imaging tasks through personalization. Specifically, our method constructs an approximation of $\mathbb{X}_{h}$ using the learnable biological prior knowledge $\mathbb{O}$, via decomposition, invariance, and equivariance constraints during pre-training (refer to Sec. 3.2). The learned approximation of $\mathbb{X}_{h}$ can then be utilized to enhance performance in downstream generalization tasks, irrespective of whether a domain gap exists between the pre-training data and downstream data.
@@
Importantly, this paper demonstrates that obtaining a personalized invariant representation, $\mathbb{X}_{h}$, is feasible through our approach, and such invariance leads to generalization improvements across various medical tasks. To validate our methodology, we conduct experiments on modality transfer and missing modality segmentation tasks, addressing not only the homogeneous generalization of MRI but also the rarely explored heterogeneous generalization, such as generalization between PET and CT. Our findings reveal that our approach successfully captures comprehensive, personalized information even when only partial modalities are available for a given individual (see Fig. 3). Moreover, extensive experiments on both homogeneous (Sec. 4) and heterogeneous (Secs. 5 and 6) generalization demonstrate that our approach can be adapted for downstream tasks and surpasses current state-of-the-art (SOTA) methods in multiple tasks, validating its superiority. We will publicly release our code, checkpoints, and data splits upon acceptance.
@@ 2 Related work
Medical generalization tasks currently concentrates on homogeneous generalization, introducing tasks, such as modality transfer and missing modality segmentation for structural modalities — Flair, T1, T2, and T1ce of MRI- in brain tumor segmentation [53], or between MRI and CT [51] for modality transfer. [34] propose an approach for heterogeneous generalization in terms of modality transfer, but only tailored for transferring PET to CT. In terms of tasks, most current studies focus on one specific generalization task, either segmentation [5, 43, 12, 52, 44, 52] or modality transfer [22, 54, 13, 35, 26, 30, 38, 11, 33, 25, 47, 51]. This paper aims to develop an approach that is feasible for different downstream tasks under both homogeneous and heterogeneous modality generalization. Our approach aims to learn the $\mathbb{X}_{h}$ through pre-training; we list related medical pre-training work [40, 46, 8, 23] here. A notable work among them is [23], which extracts class-specific anatomical invariance. However, they only focus on a single modality. Such single-modality approaches may not be able to construct $\mathbb{X}_{h}$ for improving the generalization across modalities. The generalization in medical images also connects to alignment in multi-domain generalization for natural images [14, 28, 27, 20, 39]. Please refer to Appendix A for a detailed literature review.
@@ 3 Learning personalized invariant representation for medical generalization
Preliminaries. In this paper, we denote the encoder as $\mathcal{E}$ and its corresponding decoder as $\mathcal{D}$. For an individual human being $h\in\mathcal{H}$, the corresponding medical images are represented as $X_{h}={X^{i}_{h},X^{j}_{h},\dots,X^{k}_{h}}$, where $i,j,\dots,k\in\mathcal{M}$, and $\mathcal{M}$ represents the set of all possible modality combinations. We denote the intermediate features produced by $\mathcal{E}(X_{h})$ and $\mathcal{E}(X^{i}_{h})$ as $x_{h}$ and $x^{i}_{h}$, respectively. The final layer features from the encoder are represented as $z_{h}$ and $z^{i}_{h}$. The learned approximation of $\mathbb{X}_{h}$ is denoted as ${\mathbb{X}_{h}}^{\prime}$. Additionally, we define the geometric warping function $\phi^{i}\in\Phi$, where $\phi^{i}(X^{i}_{h})\in\mathcal{X}$, and $\Phi$ denotes the set of all possible geometric warping functions. Finally, $I(\cdot;\cdot)$ and $P(\cdot)$ represent mutual information and probability distribution, respectively.
Before addressing the problem for both homogeneous and heterogeneous generalization, we introduce the $\mathbb{X}_{h}$ Hypothesis for medical imaging:
@@
Consider the set $\mathcal{M}$ of all possible modality combinations and the set $\Phi$ of all possible geometric transformations (e.g., SO(3)) transformations, for example, rotations corresponding to different poses of the person. There exists a personalized invariant representation $\mathbb{X}_{h}$ for an individual from the population $h\in\mathcal{H}$, which can be decomposed into modality-specific images $X^{i}_{h}$ given a modality combination $i\in\mathcal{M}$ as a condition:
@@
Figure 2: Left: Overall framework of learning $\mathbb{X}_{h}$. Right: Diagrams of differences between previous methods of learning $\mathbb{Z}_{h}$ and our proposed method of learning $\mathbb{X}_{h}$.
@@
3.1 Using global prior for better personalized invariant representation
To learn a better approximation of $\mathbb{X}_{h}$, we leverage a global biological prior, denoted as $\mathbb{O}$. If $\mathbb{O}$ can be learned, representations from any modality can complete themselves by retrieving the missing knowledge from $\mathbb{O}$, forming a better approximation of $\mathbb{X}_{h}$. Empirically, we initialize a learnable tensor as $\mathbb{O}$. As shown in Fig. 1, the representation $z^{i}_{h}$ retrieves its missing knowledge from $\mathbb{O}$ via attention: ${z^{i}_{h}}^{\prime}:=attn(query:z^{i}_{h},key:\mathbb{O},value:\mathbb{O})$. The original representation and the retrieved knowledge are then fused through convolution: $\mathbb{X}^{i}_{h}:=conv({z^{i}_{h}}^{\prime},z^{i}_{h})$. If the model is well-trained under the constraints of equivariance, invariance, and decomposition, the fused feature $\mathbb{X}^{i}_{h}$ becomes $\mathbb{X}^{\prime}_{h}$, a good approximation of $\mathbb{X}_{h}$. The details of these constraints are discussed in Sec. 3.2.
3.2 Constraints of equivariance, invariance, and decomposition
@@
$\min_{\mathcal{D},\mathcal{F}}Dis(\phi^{i},\mathcal{F}(z^{i}_{h})),$ (4)
where $\mathcal{F}:\mathcal{F}(z^{i}_{h})\to{\phi^{i}}^{\prime}$ extracts the geometric transformation and $Dis(\dot{,}\dot{)}$ denotes the distance measurement between ${\phi^{i}}^{\prime}$ and $\phi^{i}$. Empirically, following [40], we also adopt rotation as the geometric transformation, predicting the angle categories of input sub-volume is rotated. Under this case, $\Phi_{R}$ is defined as rotations at $[0,90,180,270]$ degrees along the z-axis, and $\phi_{r}^{i}\in\Phi_{R}$ is the ground truth rotation categories. $\mathcal{F}{z^{i}_{h}}$ produces the softmax probabilities of rotation categories. The loss is in the form of:
@@
Task T1→T2 T2 → Flair T1 → T1ce
Method PSNR↑ NMSE↓ SSIM↑ PSNR↑ NMSE↓ SSIM↑ PSNR↑ SSIM↑
2D
Pix2Pix [22] 24.624 0.109 0.874 24.82 0.0250 0.846 27.05 0.858
CycleGAN [54] 23.535 0.155 0.837 23.418 0.164 0.825 30.13 0.906
NICEGAN [7] 23.721 0.148 0.840 23.643 0.148 0.829 - -
GcGAN [13] - - - 29.98 - 0.917 25.98 0.872
CUT [35] - - - 23.54 - 0.819 26.27 0.846
RegGAN [26] 24.884 0.094 0.881 24.576 0.112 0.852 31.36 0.930
ResViT [10] 25.578 0.088 0.895 24.825 0.108 0.861 31.46 0.932
Diffusion [11] - - - 31.98 - 0.930 29.22 0.921
MD-Diff [47] - - - 30.76 - 0.934 33.08 0.948
3D
Pix2Pix 23.740 0.138 0.835 23.508 0.152 0.822 - -
CycleGAN 25.181 0.097 0.887 24.602 0.113 0.854 - -
EaGAN [50] 24.884 0.094 0.881 24.576 0.112 0.852 - -
MS-SPADE [25] 25.818 0.079 0.904 25.074 0.098 0.867 26.119 0.912
Ours 30.756 0.065 0.944 32.224 0.046 0.941 34.547 0.955
Table 1: Modality transfer results of MRI on BRATS23: Comparison between previous methods and our method. The best results are highlighted in blue. Results denoted with are gained from [25], while the results denoted with are gathered from [47].
Target T1 T1ce T2 Flair
T1 Ours 43.472 0.003 0.996 34.547 0.045 0.955 30.756 0.065 0.944 31.693 0.049 0.937
T1ce Ours 34.077 0.020 0.962 46.663 0.003 0.996 30.775 0.063 0.942 32.224 0.046 0.941
T2 Ours 32.646 0.028 0.955 33.857 0.051 0.949 43.653 0.006 0.991 32.224 0.046 0.941
Flair Ours 32.752 0.026 0.951 33.471 0.055 0.944 30.571 0.068 0.940 43.624 0.004 0.995
Table 2: Modality transfer results of MRI on BRATS23: Comparison between the previous method and ours for transfer between all four modalities. The averaged results of metrics for all validation samples are listed. The best results are highlighted in blue. Please refer to Appendix 9 for the standard derivations according to each result.
@@
$\mathcal{L}_{decom}\!=\!\sum_{1}^{|\mathcal{M}|}\left\|{\phi^{i}}^{-1}\!\left(\mathcal{D}\left({\mathbb{X}_{h}}^{\prime}|\phi^{i}(x^{i}_{h})\right)\right),X_{h}\right\|^{2},$ (6)

### v4

# Towards a Universal 3D Medical Multi-modality Generalization via Learning Personalized Invariant Representation

@@ Abstract
Variations in medical imaging modalities and individual anatomical differences pose challenges to cross-modality generalization in multi-modal tasks. Existing methods often concentrate exclusively on common anatomical patterns, thereby neglecting individual differences and consequently limiting their generalization performance. This paper emphasizes the critical role of learning individual-level invariance, i.e., personalized representation $\mathbb{X}_{h}$, to enhance multi-modality generalization under both homogeneous and heterogeneous settings. It reveals that mappings from individual biological profile to different medical modalities remain static across the population, which is implied in the personalization process. We propose a two-stage approach: pre-training with invariant representation $\mathbb{X}_{h}$ for personalization, then fine-tuning for diverse downstream tasks. We provide both theoretical and empirical evidence demonstrating the feasibility and advantages of personalization, showing that our approach yields greater generalizability and transferability across diverse multi-modal medical tasks compared to methods lacking personalization. Extensive experiments further validate that our approach significantly enhances performance in various generalization scenarios.
@@ 1 Introduction
Figure 1: A diagram of medical modalities and individual differences. These variations are significant and deserve further study by the medical intelligence community.
Three-dimensional radiological medical images, generated through specialized techniques and radiopharmaceuticals, excel at highlighting specific physiological features, collectively providing a comprehensive view of a patient’s structural and functional characteristics. As illustrated in Fig. 1, current research in medical intelligence mainly targets structural modalities like Magnetic Resonance Imaging (MRI) [58] and Computed Tomography (CT) [36, 56] scans. Other studies [54] focus on the functional modalities associated with biochemistry, like Positron Emission Tomography (PET) scans. Notably, all modalities exhibit significant individual variations. Individuals differ fundamentally from population averages [50], making anatomical and metabolic variations evident. In clinical settings, financial or physical limitations may restrict access to certain modalities, highlighting the need for generalization in medical image analysis across different modalities and individuals. The generalization faces significant challenges due to modality and individual variations. We classify generalization tasks into two types: homogeneous generalization within structural or functional modalities, e.g., MRI sequences Fig. 1; and heterogeneous generalization across both structural and functional modalities, such as CT and PET scans.
A well-generalized medical model should integrate insights from all available modalities to support downstream tasks, relying on condition invariance [10, 43]. Some methods enhance generalization through class-level anatomy invariance [25], which may not fit models based on functional biochemistry and may degrade transferability for unseen individual domains. Single-modal task transfer research [44, 51, 25] might not translate to multi-modal scenarios. Homogeneous generalization in medical tasks often employs cross-modality transfer [32, 27, 56] or tackles missing modalities [34, 7, 40, 41, 56] in MRI or CT, while heterogeneous generalization is insufficiently explored. Furthermore, these limited efforts often focus on modality or class-level invariance, neglecting individual differences and impeding a model’s capacity to generalize effectively across patient populations and modalities.
Figure 2: A diagram of patient domains and medical modalities linked to our method. In the Pre-training stage, static mappings between domains and modalities are implicitly learned. The fine-tuning stage addresses gaps between seen and unseen patients.
This paper reveals that both homogeneous and heterogeneous settings for multi-modality generalization can be tackled within the framework of personalization, i.e., learning the personal-level invariant representation. Based on medical imaging principles, mappings from an individual’s identical biological profile to various modalities remain static across individuals. Learning personal-level invariant representations, which apply across multiple modalities, enhances cross-modality generalization and transferability to novel individuals. To formalize this approach, the hypothesis of a personalized invariant representation is introduced, denoted as $\mathbb{X}_{h}$, which exists for each individual in multi-modal generalization. Hypothesis  3.1 rigorously details the constraints and properties of $\mathbb{X}_{h}$.
Building on this hypothesis, a general approach termed PUIR is proposed to augment the generalization of various medical imaging tasks through personalization. As shown in Fig. 2, our approach consists of two stages: (1) The pre-training stage constructs an approximation of $\mathbb{X}_{h}$ using the learnable biological prior knowledge $\mathbb{O}$, via decomposition, invariance, and equivariance constraints during pre-training (refer to Sec. 3.1.1), focusing on individual-level modality generalization. (2) The fine-tuning stage, employs standard downstream training methods to adapt the modality encoders derived from the first stage to mitigate the gap between seen and unseen individuals. A key distinction between our method and previous pre-training strategies lies in our specific focus on enhancing generalization across various medical modalities and individuals.
@@
Importantly, this paper theoretically demonstrates that obtaining a personalized invariant representation, $\mathbb{X}_{h}$, is feasible through our approach, and such invariance leads to generalization improvements across various medical tasks (see Sec. 3.3). To validate our methodology, we conduct experiments on modality transfer and missing modality segmentation tasks, addressing not only the homogeneous generalization of MRI but also the rarely explored heterogeneous generalization, such as generalization between PET and CT. Our findings reveal that our approach successfully captures comprehensive, personalized information even when only partial modalities are available for a given individual. Moreover, extensive experiments on both homogeneous (Sec. 4.1) and heterogeneous (Secs. 4.2 and 4.3) generalization demonstrate that our approach can be adapted for downstream tasks and surpasses current state-of-the-art (SOTA) methods in multiple tasks, validating its superiority. Our code and data splits are available at https://github.com/zhaorui-tan/PUIR_ICCV25.
@@ 2 Related work
Medical generalization tasks currently concentrates on homogeneous generalization, introducing tasks, such as modality transfer and missing modality segmentation for structural modalities — Flair, T1, T2, and T1ce of MRI- in brain tumor segmentation [58], or between MRI and CT [56] for modality transfer. [38] proposes an approach for heterogeneous generalization in terms of modality transfer, but only tailored for transferring PET to CT. In terms of tasks, most current studies focus on one specific generalization task, either segmentation [6, 47, 14, 57, 48, 33] or modality transfer [24, 60, 15, 39, 28, 32, 42, 13, 36, 27, 52, 56]. This paper aims to develop an approach that is feasible for different downstream tasks under both homogeneous and heterogeneous modality generalization. Our approach aims to learn the $\mathbb{X}_{h}$ through pre-training; we list related medical pre-training work [44, 51, 9, 25, 37] here. A notable work among them is [25], which extracts class-specific anatomical invariance. However, they only focus on a single modality. Such single-modality approaches may not be able to construct $\mathbb{X}_{h}$ for improving the generalization across modalities. The generalization in medical images also connects to alignment in multi-domain generalization for natural images [16, 30, 29, 22, 43]. Please refer to Appendix B for a detailed literature review.
Figure 3: Left: Overall framework of learning $\mathbb{X}_{h}$ during the pre-training stage. Right: Diagrams of differences between previous methods of learning $\mathbb{Z}_{h}$ and our proposed method of learning $\mathbb{X}_{h}$.
@@ 3 Learning personalized invariant representation for medical generalization
Notations. The encoder is denoted to $\mathcal{E}$, while its associated decoder is defined as $\mathcal{D}$. An individual human entity, denoted as $h\in\mathcal{H}$, with corresponding medical images $X_{h}={X^{i}_{h},X^{j}_{h},\dots,X^{k}_{h}}$, whereby $i,j,\dots,k\in\mathcal{M}$, and $\mathcal{M}$ cover all conceivable modality combinations. Intermediate features generated by $\mathcal{E}(X_{h})$ and $\mathcal{E}(X^{i}_{h})$ are designated as $x_{h}$ and $x^{i}_{h}$, respectively. The encoder’s final layer features are $z_{h}$ and $z^{i}_{h}$. The learned approximation of $\mathbb{X}_{h}$ is expressed as ${\mathbb{X}_{h}}^{\prime}$. The geometric warping function is $\phi^{i}\in\Phi$, with $\phi^{i}(X^{i}_{h})\in\mathcal{X}$ and $\Phi$ as the ensemble of all possible warping functions. Conclusively, $I(\cdot;\cdot)$ and $P(\cdot)$ signify mutual information and probability distribution, respectively.
Personalized invariant representation hypothesis. Before addressing the problem for both homogeneous and heterogeneous generalization, we introduce the personalized invariant representation hypothesis, termed as $\mathbb{X}_{h}$ Hypothesis, for medical imaging:
@@
Consider the set $\mathcal{M}$ of all possible modality combinations and the set $\Phi$ of possible geometric transformations (e.g., SO(3)), such as rotations corresponding to different postures of an individual. For a given individual from the population $h\in\mathcal{H}$, there exists a personalized invariant representation $\mathbb{X}_{h}$, which can be decomposed into modality-specific images $X^{i}_{h}$ conditioned on a modality combination $i\in\mathcal{M}$. This relationship is formally expressed as:
@@
3.1 Pre-training stage
To learn a better approximation of $\mathbb{X}_{h}$, we leverage a learnable biological prior, denoted as $\mathbb{O}$. If $\mathbb{O}$ can be learned, representations from any modality can complete themselves by retrieving the missing knowledge from $\mathbb{O}$, forming a better approximation of $\mathbb{X}_{h}$. Empirically, we initialize a learnable tensor as $\mathbb{O}$. As shown in Fig. 3, the representation $z^{i}_{h}$ retrieves its missing knowledge from $\mathbb{O}$ via attention: ${z^{i}_{h}}^{\prime}:=attn(query:z^{i}_{h},key:\mathbb{O},value:\mathbb{O})$. The original representation and the retrieved knowledge are then fused through convolution: $\mathbb{X}^{i}_{h}:=conv({z^{i}_{h}}^{\prime},z^{i}_{h})$. If the model is well-trained under the constraints of equivariance, invariance, and decomposition, the fused feature $\mathbb{X}^{i}_{h}$ becomes $\mathbb{X}^{\prime}_{h}$, a good approximation of $\mathbb{X}_{h}$. The details of these constraints are discussed in Sec. 3.1.1.
3.1.1 Constraints of equivariance, invariance, and decomposition
@@
$\min\nolimits_{\mathcal{D},\mathcal{F}}Dis(\phi^{i},\mathcal{F}(z^{i}_{h})),$ (4)
where $\mathcal{F}:\mathcal{F}(z^{i}_{h})\to{\phi^{i}}^{\prime}$ extracts the geometric transformation and $Dis(\dot{,}\dot{)}$ denotes the distance measurement between ${\phi^{i}}^{\prime}$ and $\phi^{i}$. Empirically, following [44], we also adopt rotation as the geometric transformation, predicting the angle categories of input sub-volume is rotated. Under this case, $\Phi_{R}$ is defined as rotations at $[0,90,180,270]$ degrees along the z-axis, and $\phi_{r}^{i}\in\Phi_{R}$ is the ground truth rotation categories. $\mathcal{F}{z^{i}_{h}}$ produces the softmax probabilities of rotation categories, predicting which kind of rotation is applied, and loss is in the form of:
@@
$\mathcal{L}_{decom}\!=\!\sum\nolimits_{1}^{|\mathcal{M}|}\left\|{\phi^{i}}^{-1}\!\left(\mathcal{D}\left({\mathbb{X}_{h}}^{\prime}|\phi^{i}(x^{i}_{h})\right)\right),X_{h}\right\|^{2},$ (6)
@@
Task T1→T2
Method PSNR↑ NMSE↓ SSIM↑
2D
Pix2Pix [24] 24.624 0.109 0.874
CycleGAN [60] 23.535 0.155 0.837
NICEGAN [8] 23.721 0.148 0.840
RegGAN [28] 24.884 0.094 0.881
ResViT [12] 25.578 0.088 0.895
3D
Pix2Pix 23.740 0.138 0.835
CycleGAN 25.181 0.097 0.887
EaGAN [55] 24.884 0.094 0.881
Ours (PUIR) 30.756 0.065 0.944
Task T2 → Flair
Method PSNR↑ NMSE↓ SSIM↑
2D
Pix2Pix [24] 24.82 0.0250 0.846
CycleGAN [60] 23.418 0.164 0.825
NICEGAN [8] 23.643 0.148 0.829
RegGAN [28] 24.576 0.112 0.852
ResViT [12] 24.825 0.108 0.861
Diffusion [13] 31.98 - 0.930
MD-Diff [52] 30.76 - 0.934
3D
Pix2Pix 23.508 0.152 0.822
CycleGAN 24.602 0.113 0.854
EaGAN [55] 24.576 0.112 0.852
Ours (PUIR) 32.224 0.046 0.941
Task T1 → T1ce
Method PSNR↑ SSIM↑
2D
Pix2Pix [24] 27.05 0.858
CycleGAN [60] 30.13 0.906
GcGAN [15] 25.98 0.872
CUT [39] 26.27 0.846
RegGAN [28] 31.36 0.930
ResViT [12] 31.46 0.932
Diffusion [13] 29.22 0.921
MD-Diff [52] 33.08 0.948
3D
Ours (PUIR) 34.547 0.955
Table 1: Modality transfer results of MRI on BRATS23: Comparison between previous one-to-one modality transfer methods and our method. The best results are highlighted in blue. Results denoted with are gained from [27]; with are gathered from [52].